OpenStudy (anonymous):

Find the equation of the line passing through points (5, 7) and (0, 2). Write the equation in slope-intercept form.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(5\quad ,&7)\quad &(0\quad ,&2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug and chug}\)